The copper grade of a matte is 30%. Find the weight of Cu2S and FeS in 1 ton of the matte. If
the same matte is charge in a converter find the volume of air during slag-off phase and blister
formation phase of conversion.
1
Expert's answer
2020-11-23T06:57:28-0500

PART I

Copper Matte

mCu2S.nFeS

RMM of Cu2S=160g/mol

RMM of FeS=88g/mol

0.3= [160m×128/160]/[160m+88n]

48m+26.4n=128m

128m-48m=26.4n

m/n=33/100

Formula of the compound

33Cu2S.100FeS

RMM of the compound

=5280+8800

=14080g

Mass of Cu2S from 1ton= [5280g/14080]100%

                                            =375000g

                                               =375kg

Mass of FeS from 1 ton= [1000000-375000]/1000

                                      =625kg

PART II

The amount of the flux= 75%of 100kg

                                     = [75/100]1000

                                       =750kg

Weight of SiO2 slag

Let X be the weight of slag SiO2

75% [750] =0.32x

X=1757.81kg

Let Y be the weight of FeS oxidized in one blow Fe balance:

(56/88) ×Y + 5%× 750× (56/120) = 59% ×1757.81 × (56/72) Y = 1241.65kg=14.11kg/mol

Oxygen required for 1 blow:

FeS + 3/2O2 = FeO + SO2

Oxygen required=14.11[3/2] =21.16kk/mol

FeS2 + 5/2O2 = FeO + 2SO2

Oxygen= [5/100] × [750/120] × [5/2] = 0.3125kg/mol

Total O2 =0.3125kg/mol+21.16kk/mol

             =21.4725kg/mol

Cu2S + O2 = Cu(l) + SO2

Total Cu2 S = 375kg/160= 2.3438kg/mol

Oxygen needed= 2.3438kg/mol

Total oxygen needed= 2.3438kg/mol+21.4725kg/mol

                                    =23.81625kg/mol

Volume of air = [23.81625/0.21] [22.4]

                       =2540.4m3